# DynamicRAG: Leveraging Outputs of Large Language Model as Feedback for Dynamic Reranking in Retrieval-Augmented Generation

### v1

@@ 4.3 Implementation Details
Our training data comprises 150k diverse instruction-output pairs, drawn from Alpaca Taori et al. (2023), KILT Petroni et al. (2021), ASQA Stelmakh et al. (2022), and OpenBookQA Mihaylov et al. (2018). We employ three models as the base LMs for our dynamic reranker and generator: LLaMA2-7B, LLaMA2-13B, and LLaMA3-8B. For the retriever model, we use the off-the-shelf Contriever-MS MARCO Izacard et al. (2022) as the default retriever, retrieving up to 45 documents for LLaMA3 and 20 documents for LLaMA2 per input, tailored to their respective context window sizes. Unless otherwise specified, the retriever and generator share the parameters. Additional training details can be found in the Appendix B.4.
@@
For the dynamic reranker, we set the temperature to 0.2 to enhance output diversity. For the generator, we use a temperature of 0 to ensure output stability and reproducibility. By default, we use the top 45 documents from Contriever-MS MARCO Izacard et al. (2022) as input to the reranker. In contrast, all baseline methods use the top 10 documents from Contriever-MS MARCO as input to ensure a fair comparison.
@@
GPT-3.5-Turbo Ouyang et al. (2022) Unknown 77.8 82.5 85.7 52.1 56.6 62.4 69.5
GPT-4 OpenAI et al. (2024b) Unknown 79.3 83.2 85.1 53.2 57.0 61.0 69.8
@@
BM25 Robertson and Zaragoza (2009) N/A 38.0 50.7 60.1 57.5 63.0 67.5 56.1
Contriever Izacard et al. (2022) Unknown 73.6 80.2 84.8 53.1 58.7 62.4 68.8
monoT5 Nogueira et al. (2020) $\sim$ 503k 75.6 80.9 84.9 54.8 60.2 63.3 70.0
RankLLaMA Ma et al. (2024) $\sim$ 503k 77.8 83.1 86.0 57.1 62.1 64.8 71.8
ChatQA-1.5 (LLaMA3-8B) Liu et al. (2024b) N/A 68.2 75.7 82.0 37.4 45.0 53.6 60.3
RankRAG (LLaMA3-8B) Yu et al. (2024b) $\sim$ 50k 80.3 84.0 86.3 57.6 61.8 65.2 72.5
@@
GENRE Cao et al. (2021) $\sim$ 406k 61.4 - - 34.0 - - -
Re3eval Song et al. (2024) $\sim$ 240k 65.4 - - 44.2 - - -
SEAL Bevilacqua et al. (2022) Unknown 68.2 - - 51.0 - - -
@@ 4.5.1 Impact of Reranker and Generator Size
4.5.2 Effect of Key Components on DynamicRAG Performance
4.5.3 Performance with different retrievers
We conducted experiments to compare the performance of Vanilla-RAG and DynamicRAG on three different benchmarks: NQ, HotpotQA, and ASQA, using different retrievers (DPR Karpukhin et al. (2020), Contriever, MonoT5), as shown in Figure 3. It can be observed that as the retriever models improve, both approaches exhibit better performance on downstream tasks. This also demonstrates the robustness of our model.
4.6 Further Analysis
@@ 4.6.1 Efficiency
Figure 4: Comparison of different RAG models in terms of efficiency and effectiveness. Models closer to the top-left corner achieve better overall performance.
We evaluate the maximum number of LLM calls required by different RAG models to generate an answer when the retriever returns 20 documents. The x-axis represents the number of LLM calls, while the y-axis denotes the average performance on the NQ benchmark. It is evident that the closer a model is to the top-left corner, the better it performs, as both effectiveness and efficiency are optimized. Our model is positioned in the top-left corner, demonstrating superior performance compared to other models. Specifically, our model can generate answers with only two LLM calls when the retriever returns 20 documents.
@@ 4.6.2 Reranked Document Distribution
5 Related Work
5.1 Retrieval-Augmented Generation
RAG enhances LLM performance by integrating external knowledge, improving factual accuracy and contextual understanding Lewis et al. (2020); Guu et al. (2020). Studies have explored retrieval-based next-token prediction Khandelwal et al. (2020); Ram et al. (2023) and end-to-end fine-tuning for seamless integration with retrieved content Borgeaud et al. (2022); Izacard et al. (2023); Zhang et al. (2024). Combining RAG with RLHF offers promising advancements. Asai et al. (2023) further optimize RAG by introducing special tokens for adaptive retrieval and reflection, fine-tuning LLMs with an additional critic model. In related work, Ke et al. (2024) optimized RAG by training a bridge model to refine the retriever-LLM connection. While sharing similarities, our approach differs in key ways: (1) We optimize the reranker by treating it as an agent that interacts with the generator to collect training trajectories. (2) We jointly train the reranker and generator, finding that shared parameters improve adaptation to downstream tasks.
5.2 LLM for Reranking
The application of LLMs in passage reranking has gained significant attention, with existing methods broadly categorized into point-wise, pair-wise, and list-wise approaches. Point-wise methods assess query-passage relevance in isolation using relevance or query generation Liang et al. (2023); Drozdov et al. (2023); Sachan et al. (2023). Pair-wise methods compare passage pairs to establish relative relevance Qin et al. (2024). List-wise methods rank passages holistically, aligning with Learning to Rank techniques Sun et al. (2023); Ma et al. (2023); Chao et al. (2024). Recent innovations, such as zero-shot reranking with fine-tuned open-weight LLMs Pradeep et al. (2023); Liu et al. (2024a) and logit-based approaches Gangi Reddy et al. (2024); Chen et al. (2024), aim to address these issues but often require specialized fine-tuning or face scalability constraints. Our approach treats the LLM-based reranker as an agent, first using behavior cloning to mimic expert trajectories. Subsequently, it interacts with the environment (the generator) to generate trajectories for further optimization.
@@ 6 Conclusion
In this work, we proposed DynamicRAG, a novel reinforcement learning-based framework for optimizing the reranking process in RAG systems. By modeling the reranker as an RL agent and leveraging rewards derived from the quality of LLMs’ responses, our approach enables dynamic adjustment of the order and number of retrieved documents based on the query. This dynamic reranking mechanism enhances both the relevance of selected documents and the overall system efficiency. Extensive evaluations on seven knowledge-intensive datasets demonstrate that DynamicRAG consistently outperforms existing fine-tuned and prompting-based approaches, achieving state-of-the-art performance.
@@ B.3 Dynamic Reranker Module Enhances Closed-Source Model RAG Performance
Many studies Yu et al. [2024b], Chen et al. [2023] have demonstrated that increasing the number of input documents introduces more irrelevant information, ultimately degrading model performance. Therefore, dynamically adjusting $k$ remains crucial, even for strong models. In fact, since our reward mechanism leverages responses and ground-truth outputs, our approach is fully applicable to closed-source, robust models such as GPT-4o. Experimental results validating this claim are presented below:
@@
B.4 Experiment Details
B.4.1 Training Details
We train our models on 8 Nvidia A100 GPUs, each with 80GB of memory. Fine-tuning is performed for 1 epoch with an effective batch size of 4, achieved by setting a per-device batch size of 2 and accumulating gradients over 2 steps. The learning rate is configured at 1e-5, with a 10% warmup ratio and a cosine learning rate schedule. Maximum token lengths are set to 4,096 for LLaMA2 and 8,192 for LLaMA3, adhering to the training configuration. Multi-GPU distributed training is efficiently handled using DeepSpeed Stage 3, with Bfloat16 precision to optimize memory usage. FlashAttention enhances the efficiency of long-context training. For behavior cloning, we employ MonoT5 Nogueira et al. [2020] as the expert model, set $\tau$ as 0.8 and constrain the number of documents to a maximum of 15, since many works Asai et al. [2023], Yu et al. [2024b] only use the top-10 as the input for the generator, we aim to obtain a relatively fair comparison, so we do not set k too large. For $\alpha$, $\beta$, $\gamma$, $\lambda$, and $\delta$ in reward functions, we simply use (0.2, 0.2, 0.2, 0.2, 0.2). For sampling trajectories, we use temperature to 1.0 and top p to 0.9. For DPO training, we used the following configuration: a learning rate of 5e-6, 2 epochs, a 10% warmup ratio, and a cosine learning rate schedule. The batch size was set to 1, and accumulating gradients over 4 steps. For inference, we leverage the same A100 GPUs and utilize vLLMs to accelerate inference time.
@@ B.4.1 Training Details
B.4.2 Retriever Setting
B.4.3 Training Dataset
B.4.4 Evaluation Setting
B.5 Qualitative Examples

### v2

@@ 4.3 Implementation Details
Our training data comprises 150k diverse instruction-output pairs, drawn from Alpaca [52], KILT [41], ASQA [49], and OpenBookQA [35]. We employ three models as the base LMs for our dynamic reranker and generator: LLaMA2-7B, LLaMA2-13B, and LLaMA3-8B. For the retriever model, we use the off-the-shelf Contriever-MS MARCO [16] as the default retriever, retrieving up to 45 documents for LLaMA3 and 20 documents for LLaMA2 per input, tailored to their respective context window sizes. Unless otherwise specified, the retriever and generator share the parameters. Additional training details can be found in the Appendix C.
@@
For the dynamic reranker, we set the temperature to 0.2 to enhance output diversity. For the generator, we use a temperature of 0 to ensure output stability and reproducibility. By default, we use the top 45 documents from Contriever-MS MARCO [16] as input to the reranker. In contrast, all baseline methods use the top 10 documents from Contriever-MS MARCO as input to ensure a fair comparison.
@@
GPT-3.5-Turbo [40] Unknown 77.8 82.5 85.7 52.1 56.6 62.4 69.5
GPT-4 [39] Unknown 79.3 83.2 85.1 53.2 57.0 61.0 69.8
@@
BM25 [46] N/A 38.0 50.7 60.1 57.5 63.0 67.5 56.1
Contriever [16] Unknown 73.6 80.2 84.8 53.1 58.7 62.4 68.8
monoT5 [37] $\sim$ 503k 75.6 80.9 84.9 54.8 60.2 63.3 70.0
RankLLaMA [34] $\sim$ 503k 77.8 83.1 86.0 57.1 62.1 64.8 71.8
ChatQA-1.5 (LLaMA3-8B) [32] N/A 68.2 75.7 82.0 37.4 45.0 53.6 60.3
RankRAG (LLaMA3-8B) [60] $\sim$ 50k 80.3 84.0 86.3 57.6 61.8 65.2 72.5
@@
GENRE [4] $\sim$ 406k 61.4 - - 34.0 - - -
Re3eval [48] $\sim$ 240k 65.4 - - 44.2 - - -
SEAL [2] Unknown 68.2 - - 51.0 - - -
@@ 4.5.1 Impact of Reranker and Generator Size
4.5.2 Effect of Key Components on DynamicRAG Performance
4.5.3 Performance with different retrievers
We conducted experiments to compare the performance of Vanilla-RAG and DynamicRAG on three different benchmarks: NQ, HotpotQA, and ASQA, using different retrievers (DPR [20], Contriever, MonoT5), as shown in Figure 4. It can be observed that as the retriever models improve, both approaches exhibit better performance on downstream tasks. This also demonstrates the robustness of our model.
@@ 4.5.3 Performance with different retrievers
4.6 Model Analysis
@@ 4.6.1 Efficiency
Figure 5: Comparison of different RAG models in terms of efficiency and effectiveness. The x-axis represents the number of LLM calls, while the y-axis denotes the average performance on the NQ benchmark. Models closer to the top-left corner achieve better overall performance.
We evaluate the maximum number of LLM calls required by different RAG models to generate an answer when the retriever returns 20 documents. It is evident that the closer a model is to the top-left corner, the better it performs, as both effectiveness and efficiency are optimized. Our model is positioned in the top-left corner, demonstrating superior performance compared to other models. Specifically, our model can generate answers with only two LLM calls when the retriever returns 20 documents.
@@ 4.6.2 Reranked Document Distribution
Due to space limitations, we defer several important analyses to the Appendices. Appendix B includes ablation studies on different retrievers, Top-$N$ document selection, reward function components, and training experiments with closed-source models. Appendix C provides comprehensive implementation details, qualitative examples demonstrating our dynamic reranking approach, and complete prompt templates.
4.7 Related Work
4.7.1 Retrieval-Augmented Generation
RAG boosts LLM performance by adding external knowledge, enhancing factual accuracy and context [26, 14]. Research includes retrieval-based next-token prediction [22, 45] and end-to-end fine-tuning for better integration [3, 17, 63]. Asai et al. [1] optimizes RAG using special tokens for adaptive retrieval and reflection, fine-tuning with a critic model. In related work, Ke et al. [21] optimized RAG by training a bridge model to refine the retriever-LLM connection. While sharing similarities, our approach differs in key ways: (1) We optimize the reranker by treating it as an agent that interacts with the generator to collect training trajectories. (2) We jointly train the reranker and generator, finding that shared parameters improve adaptation to downstream tasks.
4.7.2 LLM for Reranking
LLMs are increasingly used for passage reranking, with methods generally being point-wise (assessing individual relevance via relevance or query generation [27, 9, 47]), pair-wise (comparing passage pairs for relative relevance [43]), and list-wise (holistically ranking passages like Learning to Rank [51, 33, 5]). Recent advancements like zero-shot reranking with fine-tuned open-weight LLMs [42, 30] and logit-based methods [11, 7] aim to solve issues but often need specific fine-tuning or have scalability limits. Our approach treats the LLM reranker as an agent, initially using behavior cloning to imitate expert behavior, then interacting with the generator to create trajectories for further optimization.
@@ 5 Conclusion
This work introduces DynamicRAG, a new reinforcement learning framework to optimize reranking in RAG. By modeling the reranker as an RL agent and using LLM response quality as rewards, it dynamically adjusts the order and number of retrieved documents per query. This dynamic reranking mechanism enhances both the relevance of selected documents and the overall system efficiency. Extensive evaluations on seven knowledge-intensive datasets demonstrate that DynamicRAG consistently outperforms existing fine-tuned and prompting-based approaches, achieving state-of-the-art performance.
@@ B.3 Dynamic Reranker Module Enhances Closed-Source Model RAG Performance
Many studies [60, 6] have demonstrated that increasing the number of input documents introduces more irrelevant information, ultimately degrading model performance. Therefore, dynamically adjusting $k$ remains crucial, even for strong models. In fact, since our reward mechanism leverages responses and ground-truth outputs, our approach is fully applicable to closed-source, robust models such as GPT-4o. Experimental results validating this claim are presented below:
@@
Appendix C Experiment Details
C.1 Training Details
We train our models on 8 Nvidia A100 GPUs, each with 80GB of memory. Fine-tuning is performed for 1 epoch with an effective batch size of 4, achieved by setting a per-device batch size of 2 and accumulating gradients over 2 steps. The learning rate is configured at 1e-5, with a 10% warmup ratio and a cosine learning rate schedule. Maximum token lengths are set to 4,096 for LLaMA2 and 8,192 for LLaMA3, adhering to the training configuration. Multi-GPU distributed training is efficiently handled using DeepSpeed Stage 3, with Bfloat16 precision to optimize memory usage. FlashAttention enhances the efficiency of long-context training. For behavior cloning, we employ MonoT5 [37] as the expert model, set $\tau$ as 0.8 and constrain the number of documents to a maximum of 15, since many works [1, 60] only use the top-10 as the input for the generator, we aim to obtain a relatively fair comparison, so we do not set $k$ too large. For $\alpha$, $\beta$, $\gamma$, $\lambda$, and $\delta$ in reward functions, we simply use (0.2, 0.2, 0.2, 0.2, 0.2). For sampling trajectories, we use temperature to 1.0 and top p to 0.9. For DPO training, we used the following configuration: a learning rate of 5e-6, 2 epochs, a 10% warmup ratio, and a cosine learning rate schedule. The batch size was set to 1, and accumulating gradients over 4 steps. For inference, we leverage the same A100 GPUs and utilize vLLMs to accelerate inference time.
C.1.1 Retriever Setting
C.1.2 Training Dataset
C.1.3 Evaluation Setting
@@
C.2 Qualitative Examples